# A Principled Framework for Evaluating on Typologically Diverse Languages

### v1

@@ 1 Introduction
To assess whether a language model performs well across languages, ideally, one would evaluate it on all languages in the world. However, collecting high-quality data on such a scale is not feasible. Therefore, multilingual models are evaluated on a sample of the world’s languages. To ensure generalizability, such a language sample should be diverse, with varying characteristics and properties. However, existing approaches to this sampling process have not been optimal. Previous work has established that there is no clear terminology or methodological consensus for what constitutes ‘typological diversity’ (Ploeger et al., 2024). Currently, many approaches in NLP use phylogenetic heuristics to ensure ‘typological diversity’. In this paper, we show that this approximation of typological diversity through language phylogeny has severe shortcomings, and we provide a framework for systematically selecting languages based on typological distance measures. Our framework enables two sampling methods that are widely established within the field of linguistic typology. We demonstrate how our framework can be used for diverse typological language sampling, how it can help guide dataset expansions, and how it has use beyond typology.
@@ 5.2 Language Distance Calculation
  • Binarization: For using Grambank’s multistate values in the same format as the binary features, we incorporate a binarization option. We follow the authors of Grambank by dividing each multistate value into into two binary features, which are not logically dependent. [^3] https://github.com/grambank/grambank/blob/master/docs/Grambank_most_updated_sheet.tsv
@@ 5.3.1 MaxSum Diversity
5.3.2 MaxMin Diversity
@@ 6.1 Baselines
In practice, random sampling in NLP is not truly random, because data availability plays a large role. Therefore we select the languages that are most commonly used in practice previous work in NLP that claims to have ‘typologically diverse’ language selections. Ploeger et al. (2024) annotated the language selections in papers containing such claims. For our baseline, we sort these languages based on occurrence frequency, and sample the first $k$. In case of a tie in terms of occurrence frequency, selection is random.
@@ 6.2 Metrics
The previous metrics do not measure the extent to which individual typological properties are covered. This is especially relevant for variety sampling, where rare typological features should be included. Miestamo, Bakker, and Arppe (2016) defined the measure of saturation of a typological feature as “the proportion of values, out of the maximum number of possible values, found in the sample for that feature”. Because we deal with binary features only, we calculate the feature value inclusion (FVI) per feature as the percentage of languages that include the feature. We report the average over all features. FVI can be formalized as follows:
@@ 6.3 Results and Discussion
We compare our sampling methods against baselines based on previously used methods in NLP, as introduced in Section 6.1. Figure 5 shows that our methods consistently retrieve more diverse samples than all baselines. For the pairwise metrics (MDP, FVO), this difference is especially large for smaller samples; it is easier to avoid overlap when sampling fewer languages. This is in line with earlier findings from typology, which described the trade-off between coverage and independence (Miestamo, Bakker, and Arppe, 2016). While the FVI difference across methods is small for large sample sizes, we find that our methods retrieve a higher inclusion of feature values for small samples (<20) than the baselines. This is especially relevant for NLP, where multilingual hypotheses are commonly tested on a small set of ‘typologically diverse’ languages (Median = 11; Ploeger et al. 2024). Lastly, we observe that MaxSum and MaxMin consistently retrieve samples with higher feature entropy than other methods.
@@
7 Use Cases
@@
7.1 Sampling for Fairer Multilingual Evaluation
Multilingual language technology is typically only evaluated on a handful of seemingly randomly selected languages. This lack of systematic language sampling makes it difficult to assess how multilingual such technology really is. We present a use case of our framework that demonstrates how typologically diverse test language sampling affects the generalizability of conclusions.
Here, we conduct a large-scale cross-lingual sampling comparison of subword tokenization with tokenizers of popular multilingual models. For comparability between languages, the data should ideally be parallel across languages. Previous work (Ahia et al., 2023; Petrov et al., 2023) evaluates on the FLoRes-200 dataset, as this is multi-parallel. We considerably extend upon their language coverage in our analysis, by evaluating on text from the Parallel Bible Corpus (Mayer and Cysouw, 2014), which is the largest massively multi-parallel dataset in terms of language coverage. We first match language ISO codes to Glottocodes. We then select all languages that have Grambank coverage, and control for script (i.e. we filter out non-Latin scripts). This retrieves 571 languages, which is the broadest language coverage thus far in tokenization analysis. This is important, because we aim to test to what extent the sample represent the sampling frame, and potentially the sampling universe. We select the longest bible per language, sample the 2,000 most common verses and randomly select 1,000 of those for our evaluation, while retaining multi-parallelism. We analyse tokenizers of four popular language models with tokenizers publicly available on HuggingFace [^11] https://huggingface.co: multilingual BERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), GPT2 Radford et al. (2019) and multilingual E5 (Wang et al., 2024). Similar to Ahia et al. (2023), we measure the amount of segmentation through the average number of subwords per verse. [^12] We do not use the fertility measure (Rust et al., 2021), because it assumes similar word tokenizer performance across languages. We do not use ‘premiums’, the disparity of tokenization length between parallel sentences in two languages (Petrov et al., 2023), because these are by definition relative to another language; we instead want to compare across languages directly.
We compare our sampling methods (MaxSum, MaxMin) with the only baseline that is deterministic for $k=20$: convenience. From the 571 total languages, we sample 20 with each method. For each tokenizer, we compare the spread in average number of subwords per verse for all three sampling methods, and compare those with the spread for all 571 languages in our experiment. This gives us an estimate of how generalizable the sample of 20 languages is with respect to the total 500+ languages included in the experiment.
The results are in Figure 6. For all four models, we observe that the convenience baseline retrieves a considerably lower average number of subwords, with a smaller range than other methods. This implies that evaluating on the 20 most commonly included languages in ‘typologically diverse’ samples in NLP gives an overly optimistic image of general multilingual tokenizer performance. Yet, the MaxSum and MaxMin methods retrieve averages and spreads that much more closely resemble the average and spread of all 500+ languages in the experiment. This suggests that a priori typologically informed language sampling can improve the generalizability of the results, based on a language sample.
@@ 7.3 Other Distance Maximisation
Figure 7: We use MaxMin sampling to find twenty geographically distant language coordinate pairs.
@@ 9 Conclusions
In this work, we systematically analyse common language sampling strategies in NLP, and find that these are insufficient for typologically diverse language sampling. Guided by research in linguistic typology, we propose two sampling algorithms. We compare the samples obtained by our methods with strong baselines. These samples are evaluated with four typological diversity metrics that show that our method consistently retrieves language samples with higher typological diversity. While we focus on achieving high typological distance, our framework can be used with any type of information that one wishes to make ‘diverse’ generalizable claims about. Furthermore, our method can also be used for finding typologically similar languages, e.g., with use cases in transfer learning scenarios. We recommend that future work in NLP aiming for typological generalizability use our informed selection method (while not losing sight of the human-centered perspective).
@@ Acknowledgements.
EP and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. We thank the members of LAGoM NLP group at KU Leuven, the TypNLP lab at Aalborg University and the Helsinki-NLP group at the University of Helsinki for feedback on earlier versions of this paper. We thank Robert Östling for pointing us to Anna Sjöberg’s thesis. We thank Kaius Sinnemäki and Anna Sjöberg for discussing the scope of this project with us from a typological perspective. We thank Hedvig Skirgård for helpful feedback on cropping the dataset according to coverage, removing macrolanguages, and the overall scope of this paper. Any remaining errors are our own.

### v2

@@ 1 Introduction
To assess whether a language model performs well across languages, ideally, one would evaluate it on all languages in the world. However, collecting high-quality data on such a scale is not feasible. Therefore, multilingual models are evaluated on a sample of the world’s languages. To ensure generalizability, such a language sample should be diverse, with varying characteristics and properties. However, existing approaches to this sampling process have not been optimal. Previous work has established that there is no clear terminology or methodological consensus for what constitutes ‘typological diversity’ (Ploeger et al., 2024). Currently, many approaches in NLP use phylogenetic heuristics to ensure ‘typological diversity’. In this paper, we show that this approximation of typological diversity through language phylogeny has severe shortcomings, and we provide a framework for systematically selecting languages based on typological distance measures. Our framework implements two sampling methods that are widely established within the field of linguistic typology. We demonstrate how our framework can be used for diverse typological language sampling, how it can help guide dataset expansions, and how it has use beyond typology.
@@ 2 Background
Figure 1: A diagram showing the relationship between the sampling universe, sampling frame and the sample. The objective of diverse language sampling is to draw a sample from the sampling frame that optimally reflects a larger population (frame or universe).
@@ 5.2 Language Distance Calculation
  • Binarization: For using Grambank’s multistate values in the same format as the binary features, we incorporate a binarization option. We follow the authors of Grambank by dividing each multistate value into two binary features, which are not logically dependent. [^3] https://github.com/grambank/grambank/blob/master/docs/Grambank_most_updated_sheet.tsv
@@ 5.3.2 MaxMin Diversity
Within linguistic typology, variety sampling is used to test a hypothesis ‘against the extremes’, while probability sampling is used for research questions that require more even language distributions. In Section 7, we further contrast these methods for downstream use cases in NLP, showing that MaxSum sampling is especially useful for scenarios in which a worst-case performance estimate is desired.
@@ 6.1 Baselines
In practice, random sampling in NLP is not truly random, because data availability plays a large role. Therefore we select the languages that are most commonly used in previous work in NLP that claim to have ‘typologically diverse’ language selections. Ploeger et al. (2024) annotated the language selections in papers containing such claims. For our baseline, we sort these languages based on occurrence frequency, and sample the first $k$. In case of a tie in terms of occurrence frequency, selection is random.
@@ 6.2 Metrics
The previous metrics do not measure the extent to which individual typological properties are covered. This is especially relevant for variety sampling, where rare typological features should be included. Miestamo, Bakker, and Arppe (2016) defined the measure of saturation of a typological feature as “the proportion of values, out of the maximum number of possible values, found in the sample for that feature”. As the name suggests, this metric is expected to saturate for larger language selections. Therefore, it is especially suitable for comparing smaller language samples. Because we deal with binary features only, we calculate the feature value inclusion (FVI) per feature as the percentage of languages that include the feature. We report the average over all features. FVI can be formalized as follows:
@@ 6.3 Results and Discussion
We compare our sampling methods against baselines based on previously used methods in NLP, as introduced in Section 6.1.
Figure 6 shows that our methods consistently retrieve more diverse samples than all baselines. For the pairwise metrics (MDP, FVO), this difference is especially large for smaller samples; it is easier to avoid overlap when sampling fewer languages. This is in line with earlier findings from typology, which described the trade-off between coverage and independence (Miestamo, Bakker, and Arppe, 2016). While the FVI difference across methods is small for large sample sizes, we find that our methods retrieve a higher inclusion of feature values for small samples (<20) than the baselines. This is especially relevant for NLP, where multilingual hypotheses are commonly tested on a small set of ‘typologically diverse’ languages (Median = 11; Ploeger et al. 2024). Lastly, we observe that MaxSum and MaxMin consistently retrieve samples with higher feature entropy than other methods.
@@
7 How does Typological Diversity relate to Generalizability?
In this work, we argue that increasing the linguistic representativeness of a language sample leads to better generalizability of results in multilingual NLP. Currently, multilingual language technology is typically only evaluated on a handful of seemingly randomly selected languages. This lack of systematic language sampling makes it difficult to assess how multilingually generalizable such technology really is. In this section, we assess whether sampling typologically diverse languages actually leads to better estimates of general multilingual performance. We apply our sampling framework to three downstream NLP tasks, demonstrating how sampling diverse test languages affects the generalizability of the results. Specifically, we look into natural language understanding (topic classification), natural language generation (machine translation) and subword segmentation.
7.1 Natural Language Understanding
To assess the extent to which samples drawn with different methods reflect general multilingual performance, it is important we sample from a dataset that includes sufficient typological variation. The Taxi1500 dataset (Ma et al., 2023), which spans more than 1500 languages [^10] The intersection between Grambank and Taxi1500 contains 1506 languages., covers a large part of the ‘typological space’ (see Appendix B). Importantly, its evaluation set is parallel across languages, which facilitates cross-lingual comparison. The benchmarking task of Taxi1500 is classifying the topic of bible verses, where the six possible topics are recommendation, faith, description, sin, grace and violence. These topics were manually annotated on an English bible, and applied to other languages through annotation projection to create a large-scale parallel benchmark. We compare the results of widely popular classification models (mBERT, Devlin et al., 2019; XLM-R, Conneau et al., 2020) as well as the more recently introduced Glot500-m (Imani et al., 2023).
First, in Table 2, we compare the average scores over the samples, and their similarity to the average over all languages. Ideally, diverse language sampling should lead to averages that are more comparable to the average over all languages. We observe that this is indeed the case with our sampling methods, and that especially results obtained with the convenience baseline provide overly optimistic pictures of general multilingual performance, as compared to the more realistic performance estimates obtained with our carefully sampled language selections. Importantly, we highlight differences between MaxSum and MaxMin: MaxSum gives more conservative performance estimates than MaxMin, since, similar to variety sampling, it samples the most extreme cases. This may be useful to NLP practitioners, especially in scenarios where a worst-case estimate is desired.
In Figure 7, we go beyond average results, and inspect the distribution of the results. We compare the distribution of the accuracy over all 1506 languages with samples ($k=20$) from two all methods: MaxSum, Maxmin (ours) and convenience (reflecting common practice in NLP). We observe that the convenience baseline gives consistent overestimations of multilingual performance, while the results of our systematic sampling method consistently reflect the distributions of results for all languages in the dataset. Note that, as illustrated in Figure 1, the objective of systematic language sampling is to draw a language sample from the sampling frame that optimally reflects the larger population. This implies that performance estimates on true typologically diverse languages indeed paint a more generalizable picture of multilingual model performance.
Table 2: Average results (accuracy) per sampling method ($k=20$) on Taxi1500 dataset. For the non-deterministic sampling method, we report the average over 10 runs.
Non-deterministic Methods Deterministic Methods
Rand. Rand. Fam Rand. Gen. Conv. MaxSum MaxMin All
mBERT 14.3 ±${\scriptscriptstyle\pm 1.66}$ 12.7 ±${\scriptscriptstyle\pm 1.22}$ 13.5 ±${\scriptscriptstyle\pm 1.31}$ 39.7 12.7 13.7 13.7
XLM-R-B 13.0 ±${\scriptscriptstyle\pm 2.91}$ 9.7 ±${\scriptscriptstyle\pm 1.91}$ 11.0 ±${\scriptscriptstyle\pm 2.81}$ 56.8 7.7 13.8 11.0
XLM-R-L 12.4 ±${\scriptscriptstyle\pm 3.34}$ 8.6 ±${\scriptscriptstyle\pm 2.05}$ 10.3 ±${\scriptscriptstyle\pm 3.14}$ 59.7 5.9 13.6 10.1
Glot500-m 16.8 ±${\scriptscriptstyle\pm 3.64}$ 10.8 ±${\scriptscriptstyle\pm 1.70}$ 12.9 ±${\scriptscriptstyle\pm 3.01}$ 54.6 9.6 13.9 14.4
Figure 7: Accuracy distributions of language samples ($k=20$) on the Taxi1500 evaluation set, contrasting all four deterministic sampling methods. Vertical lines represent averages per sampling method. Convenience sampling consistently over-estimates performance, as compared to MaxSum and MaxMin.
7.2 Natural Language Generation
Beyond natural language understanding, we evaluate the effect of diverse sampling on a generative task, namely machine translation (MT). We evaluate the effect of the source language on MT performance to English, using the multilingual NLLB-200 model (3.3B parameters).111111huggingface/facebook/nllb-200-3.3B The results are retrieved from the OPUS Dashboard (Tiedemann and de Gibert, 2023). Due to cross-lingual differences, MT performance can only be compared on the same target language and the same test sets. We choose to focus on investigating translation to English specifically, as this is the only language for which the dashboard contains results from all source languages in FloRes-200. English-centric MT has widely been prioritized in MT research, largely motivated by speaker numbers and information access, as argued by Bugliarello et al. (2020), among others. Furthermore, to keep the focus on typological characteristics and avoid the confounding factor of writing systems, we only take into account languages that are written in Latin script. The sampling frame contains 89 source languages in total.
Table 3: Average results (BLEU) per sampling method ($k=20$) on FloRes-200 dataset. Non-deterministic sampling methods are run 10 times.
Non-deterministic Methods Deterministic Methods
Rand. Rand. Fam Rand. Gen. Conv. MaxSum MaxMin All
34.74±${\scriptscriptstyle\pm 3.29}$ 32.74 ±${\scriptscriptstyle\pm 1.18}$ 33.00 ±${\scriptscriptstyle\pm 0.98}$ 46.64 30.05 31.52 32.69
Figure 8: Distribution of results (BLEU) of many-to-English MT with NLLB-3.3B, on the languages in Latin script from the FloRes-200 dataset. Vertical lines represent averages.
The results, as shown in Table 3 and Figure 8, demonstrate that convenience sampling gives an overestimation of performance of many-to-English MT. While performance may also be determined by the similarity to English, we see that MaxMin and MaxSum retrieve averages and distributions that resemble the distribution over the entire sampling frame. Furthermore, sampling with the MaxSum objective leads to a more conservative performance estimate than the MaxMin objective, consistently highlighting its potential for application in worst-case scenarios.
7.3 Subword Segmentation
Here, we conduct a large-scale cross-lingual sampling comparison of subword tokenization with tokenizers of popular multilingual models. For comparability between languages, the data should ideally be parallel across languages. Previous work (Ahia et al., 2023; Petrov et al., 2023) evaluates on the FLoRes-200 dataset, as this is multi-parallel. We considerably extend upon their language coverage in our analysis, by evaluating on text from the Parallel Bible Corpus (Mayer and Cysouw, 2014), which is the largest massively multi-parallel dataset in terms of language coverage. We first match language ISO codes to Glottocodes. We then select all languages that have Grambank coverage, and control for script (i.e. we filter out non-Latin scripts). This retrieves 571 languages, which is the broadest language coverage thus far in tokenization analysis. This is important, because we aim to test to what extent the sample represent the sampling frame, and potentially the sampling universe. We select the longest bible per language, and sample 1,000 commonly available verses for our evaluation, to retain multi-parallelism. We analyse tokenizers of four popular language models with tokenizers publicly available on HuggingFace [^12] https://huggingface.co: multilingual BERT (Devlin et al., 2019), LLama 3.3 70B Instruct (Dubey et al., 2024), Qwen2.5-32B (Team, 2024) and Phi 4 (Abdin et al., 2024). Similar to Ahia et al. (2023), we measure the amount of segmentation through the average number of subwords per verse. [^13] We do not use the fertility measure (Rust et al., 2021), because it assumes similar word tokenizer performance across languages. We do not use ‘premiums’, the disparity of tokenization length between parallel sentences in two languages (Petrov et al., 2023), because these are by definition relative to another language; we instead want to compare across languages directly.
We compare our sampling methods (MaxSum, MaxMin) with the only baseline that is deterministic for $k=20$: convenience. From the 571 total languages, we sample 20 with each method. For each tokenizer, we compare the spread in average number of subwords per verse for all three sampling methods, and compare those with the spread for all 570 languages in our experiment. This gives us an estimate of how generalizable the sample of 20 languages is with respect to the total 500+ languages included in the experiment.
The results are in Figure 9. For all four models, we observe that the convenience baseline retrieves a considerably lower average number of subwords, with a generally smaller range than other methods. This implies that evaluating on the 20 most commonly included languages in ‘typologically diverse’ samples in NLP gives an overly optimistic image of general multilingual tokenizer performance. Yet, MaxMin retrieves averages and spreads that more closely resemble the average and spread of all 500+ languages in the experiment. This suggests that a priori typologically informed language sampling can improve the generalizability of the results, based on a language sample.
8 Further Use Cases
@@ 8.2 Other Distance Maximisation
Figure 10: We use MaxMin sampling to find twenty geographically distant languages, based on coordinate pairs.
@@ 10 Conclusions
In this work, we systematically analyse common language sampling strategies in NLP, and find that these are insufficient for typologically diverse language sampling. Guided by research in linguistic typology, we propose two sampling algorithms. We compare the samples obtained by our methods with strong baselines. These samples are evaluated with four typological diversity metrics that show that our method consistently retrieves language samples with higher typological diversity. For downstream tasks, we show that our sampling algorithms provide a much more realistic estimate of generalizability as compared to, e.g., convenience sampling, thereby highlighting the importance of addressing typological diversity in the context of empirical NLP work. While we focus on achieving high typological distance, our framework can be used with any type of information that one wishes to make ‘diverse’ generalizable claims about. Furthermore, our method can also be used for finding typologically similar languages, e.g., with use cases in transfer learning scenarios. We recommend that future work in NLP aiming for typological generalizability use our informed selection method (while not losing sight of the human-centered perspective).
@@
Diversity of downstream sampling frames.
Figure 12: PCA plot illustrating the extent to which Taxi1500 covers Grambank’s ‘typological space’.
Figure 13: PCA plot illustrating the extent to which FloRes-200 covers Grambank’s ‘typological space’.
[…]
@@ Acknowledgements.
EP and JB are funded by the Carlsberg Foundation, under the Semper Ardens: Accelerate programme (project nr. CF21-0454). WP is funded by a KU Leuven Bijzonder Onderzoeksfonds C1 project with reference C14/23/096. This work received support from the CA21167 COST action UniDive, funded by COST (European Cooperation in Science and Technology). We thank the members of LAGoM NLP group at KU Leuven, the AAU-NLP lab at Aalborg University and the Helsinki-NLP group at the University of Helsinki for feedback on earlier versions of this paper. We thank Robert Östling for pointing us to Anna Sjöberg’s thesis. We thank Kaius Sinnemäki and Anna Sjöberg for discussing the scope of this project with us from a typological perspective. We thank Hedvig Skirgård for helpful feedback on cropping the dataset according to coverage, removing macrolanguages, and the overall scope of this paper. Any remaining errors are our own.